OpenStudy (anonymous):

solve for x: 2x-(x-3)/(8)=(1)/(2)+(x+5)/(2) PLEASE HELP :)

OpenStudy (anonymous):

x = 21/11

OpenStudy (anonymous):

HOW? show me some work please

OpenStudy (anonymous):

First, multiply both sides by 8 and you will get: 15x +3 = 4x +24 Then move the terms around and isolate x.

OpenStudy (anonymous):

you mean 16x+3?

OpenStudy (anonymous):

No, I meant 15x. 16x - x = 15x.

OpenStudy (anonymous):

ohhh